OpenStudy (anonymous):

The dimensions of a rectangle are such that its length is 3 inches more that its width. If the length and width were decreased by 1 inch, the area would be increased by 150inch sq. what are the length and width of the rectangle?

OpenStudy (anonymous):

Geometry: Rectangles Solvers Lessons Answers archive Quiz In Depth If you need immediate math help from PAID TUTORS right now, click here. (paid link) Click here to see ALL problems on Rectangles Question 378785: please help me solve this problem: the dimensions of a rectangle are such that its length is 3 inches more than its width. if the length were doubled and if the with decreased by 1 inch the area would be increased by 50 inches^2. what are the length and width of the rectangle? Answer by mananth(10567) (Show Source): You can put this solution on YOUR website! the dimensions of a rectangle are such that its length is 3 inches more than its width. if the length were doubled and if the with decreased by 1 inch the area would be increased by 50 inches^2. what are the length and width of the rectangle? ... width be x length = x+3 .. Area = x(x+3) Area = x^2+3x ... length doubled = 2(x+3) width x-1 Area = 2(x+3)*(x-1) Area = 2(x^2+2x-3) ... 2(x^2+2x-3)-(x^2+3x)=50 2x^2+4x-6-x^2-3x=50 x^2+x-6=50 x^2+x-56=0 x^2+8x-7x-56=0 x(x+8)-7(x+8)=0 (x+8)(x-7)=0 x=7 ignore -8 width = 7 inches length = x+3 = 10 inches ...

OpenStudy (anonymous):

thanks k